OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ 0.abababab\cdot \cdot \cdot \cdot \cdot \cdot \cdot \times n \hspace{.33em}\\~\\ \normalsize \text{is an integer.} \hspace{.33em}\\~\\ \normalsize \text{a and b are single digit natural number.} \hspace{.33em}\\~\\ \normalsize \text{find the greatest 3 digit value that 'n' can take.} \hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

990 ?

OpenStudy (mathmath333):

yes , but m not sure how that came

ganeshie8 (ganeshie8):

that just came haha

ganeshie8 (ganeshie8):

(jk)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ 0.abababab\cdot \cdot \cdot \cdot \cdot \cdot \cdot \times n \hspace{.33em}\\~\\ =\dfrac{ab}{99} \times n \hspace{.33em}\\~\\ }\end{align}\) how to proceed

ganeshie8 (ganeshie8):

for that stuff to be an integer for arbitrary \(a, b\) , you need \(n\) to be a multiple of \(99\) ?

OpenStudy (mathmath333):

yes

OpenStudy (mathmath333):

i m still in doubt ,what this line means exactly \(\normalsize \text{find the greatest 3 digit value that 'n' can take.} \hspace{.33em}\\~\\\)

ganeshie8 (ganeshie8):

999 is the greatest 3 digit number, yes ?

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

we have \(n = 99k\) for some integer \(k\) \(k = 2-10\) gives you 3 digit numbers, out of which 99*10 is the greatest one

OpenStudy (mathmath333):

ok and if the question was \(\normalsize \text{find the leat 3 digit value that 'n' can take.} \hspace{.33em}\\~\\\) then it would be \(99\times 2\)

ganeshie8 (ganeshie8):

thats true only when \(\gcd(99, (ab)_{10}) =1\) : \[\dfrac{ab}{99}\times n\]

ganeshie8 (ganeshie8):

example : \[\dfrac{18}{99}\times n\] least 3 digit value for \(n\) is not \(99*2\) anymore..

OpenStudy (mathmath333):

so \(ab_{10}\) should be atleast \(50\) for this \(\normalsize \text{ greatest 3 digit value that 'n' can take.} \hspace{.33em}\\~\\\)

OpenStudy (mathmath333):

if we take \(n=2*99\)